# Heat currents in qubit systems

### v4

@@ I Introduction
There is a current interest in quantum thermodynamics Alicki (1979); Barra (2015); Kosloff (2013); Kosloff and Levy (2014); Mari and Eisert (2012); Kosloff and Levy (2019); Levy et al. (2020); Colla and Breuer (2022); Linden et al. (2010). Thermodynamics is a universal theory and must from a fundamental point of view emerge from a microscopic point of view. In this context the emergence of in particular the second law of thermodynamics. In the Clausius formulation the second law implies that heat flows from a hot reservoir to a cold reservoir and this essential property should also hold in the quantum regime. In non equilibrium quantum statistical mechanics the instantaneous heat current is subject to both statistical and quantum fluctuations; however, the mean heat current must obey the second law of thermodynamics.
@@
Whereas the case of a single qubit (two-level atomic system) driven by two heat bath is easily analysed, see e.g. Aurell and Montana (2019), and discussed in the present paper, the discussion in the literature is usually focussed on composite systems Kohler et al. (2005); Kosloff and Levy (2014). Here a recent discussion has focussed on the connection between a global master equation approach as applied in the present paper and a local master equation approach in which the interaction in the composite system is ignored in the dissipative kernels. It has been shown that in the case of a local master equation approach the apparent violation of the second law can be remedied by using a protocol of repeated interactions; this scheme accounts for a work contribution which together with the heat contribution restores the validity of the second law Barra (2015); Pereira (2018); Chiara et al. (2018); Hewgill et al. (2021). In the case of coupled quantum oscillators driven by two heat reservoirs a global approach was discussed in Levy and Kosloff (2014) resolving the issue regarding the second law. We also note recent work Soret et al. (2022) dealing with the general issues regarding the consistency of quantum master equations in the analysis of quantum thermodynamics.
In the present work, motivated by the study in Levy and Kosloff (2014), we discuss a composite system of two coupled qubits driven by two heat reservoirs. Coupled qubit systems have been studied extensively but mainly in the context of quantum entanglement del Valle (2010); Liao et al. (2011); Fischer and Breuer (2013); Duan et al. (2013); Paneru et al. (2020); Schlosshauer (2019); Aolita et al. (2015); Decordi and Vidiella-Barranco (2017); Bresque et al. (2021); Hewgill et al. (2018); Hu et al. (2018); Karimi et al. (1965); in the present paper, however, the focus is on quantum thermodynamics and the second law of thermodynamics.
@@ II.3 Redfield - Lindblad - Coherences
Then issue of coherences in the steady state and currents is subtle Li et al. (2014); Zhang and Wang (2014); Wichterich et al. (2007). In recent work Fogedby (2022) alluded to in the introduction we presented a detailed discussion of the application of field theoretical methods to open quantum systems. The field theoretical approach is based on a systematic expansion in terms of diagrams and applies a condensed matter quasi-particle approximation in order to implement the Markov limit. The field theoretical method is the basis for the present analysis.
@@ III.1.1 Model
$H=\frac{\omega_{0}}{2}\sigma^{z},$ (22)
@@ III.2.1 Model
$H=\frac{\omega_{1}}{2}\sigma^{z}_{1}+\frac{\omega_{2}}{2}\sigma^{z}_{2}+\lambda(\sigma^{+}_{1}\sigma^{-}_{2}+\sigma^{-}_{1}\sigma^{+}_{2}),$ (56)
@@ III.2.3 Density matrix
Introducing the total kernel $K=K^{A}+K^{B}$, where $K^{A}$ and $K^{B}$ are given by (109) and (114), respectively, the populations $\rho_{nn}$ are determined by the linear equations $\sum_{q}K_{pp,qq}\rho_{qq}=0$. By inspection we note incidently that $\det K=0$ allowing for a solution. In terms of the parameters $s_{n}=a_{n}+b_{n}$, $n=1\cdots 4$, where $a_{n},b_{n}$ are given (100), we find the normalised populations, i.e., $\sum_{n}\rho_{nn}=1$,
@@ III.2.4 Currents
$q^{A}_{-}=\frac{1}{2}\frac{(\alpha\beta)^{2}g_{A}(\omega_{-})g_{B}(\omega_{-})\omega_{-}}{\beta^{2}g_{A}(\omega_{-})T_{A}+\alpha^{2}g_{B}(\omega_{-})T_{B}}[T_{A}-T_{B}],$ (148)
@@
$q^{A}_{-}=\frac{(\alpha\beta)^{2}g_{A}(\omega_{-})g_{B}(\omega_{-})\omega_{-}}{\beta^{2}g_{A}(\omega_{-})+\alpha^{2}g_{B}(\omega_{-})}[e^{-\omega_{-}/T_{A}}-e^{-\omega_{-}/T_{B}}],$ (152)
@@ III.3 Redfield approach
In Fig. 5 we have depicted the populations $\rho_{nn}$ as a function of the mean temperature $T=(T_{A}+T_{B})/2$ in the range $\Delta T<T<8$ with non equilibrium bias $\Delta T=(T_{A}-T_{B})/2=5.0$ with parameter choice $\omega_{1}=1.0$, $\omega_{2}=2.0$, $g_{A}(\omega_{\pm})=g_{b}(\omega_{\pm})=1.0$, and $\lambda=0.5$. The plot shows that $\rho_{44}$ and $\rho_{44}$ assumes negative value in the range $0<T\lesssim 6.0$, indicating that the Redfield approach is inadequate in analysing the coupled qubit case.
@@ IV Discussion and summary
In this paper we have discussed heat currents in two qubit systems driven by two heat reservoirs maintained at different temperatures in order to test the validity of the second law of thermodynamics in the microscopic domain, an important issue in quantum thermodynamics. The results for a single qubit system driven by two reservoirs are given in Sec. III.1 and are in accordance with the second law. The case of coupled qubits is more involved, the results are presented in Sec. III.2 and are, likewise, in agreement with the second law. We note that the current expressions in the single qubit and coupled qubit cases have a similar structure. In deriving these results it is essential to apply a global approach and operate in the basis of the full system Hamiltonian. Moreover, in order to avoid an erroneous result it is also essential to use the Lindblad master equation approach and in this way ensure a proper quantum channel with positive populations.

### v5

@@ I Introduction
There is a current interest in quantum thermodynamics Alicki (1979); Barra (2015); Kosloff (2013); Kosloff and Levy (2014); Mari and Eisert (2012); Kosloff and Levy (2019); Levy et al. (2020); Colla and Breuer (2022); Linden et al. (2010). Thermodynamics is a universal theory and must from a fundamental point of view emerge from a microscopic point of view. In this context the emergence of in particular the second law of thermodynamics. In the Clausius formulation the second law implies that heat flows from a hot reservoir to a cold reservoir and this essential property should also hold in the quantum regime. In non equilibrium quantum statistical mechanics the instantaneous heat current is subject to both statistical and quantum fluctuations, however, the mean heat current must obey the second law of thermodynamics.
@@
Whereas the case of a single qubit (two-level atomic system) driven by two heat bath is easily analysed Aurell and Montana (2019) and discussed in the present paper, the discussion in the literature is usually focussed on composite systems Kohler et al. (2005); Kosloff and Levy (2014). Here a recent discussion has focussed on the connection between a global master equation approach as applied in the present paper and a local master equation approach in which the interactions in the composite system are ignored in the dissipative kernels. It has been shown that in the case of a local master equation approach the apparent violation of the second law can be remedied by using a protocol of repeated interactions; this scheme accounts for a work contribution which together with the heat contribution restores the validity of the second law Barra (2015); Pereira (2018); Chiara et al. (2018); Hewgill et al. (2021). In the case of coupled quantum oscillators driven by two heat reservoirs a global approach was discussed in Levy and Kosloff (2014) resolving the issue regarding the second law. We also note recent work Soret et al. (2022) dealing with the general issues regarding the consistency of quantum master equations in the analysis of quantum thermodynamics.
In the present work, motivated by the study in Levy and Kosloff (2014), we discuss a composite system of two coupled qubits driven by two heat reservoirs. Coupled qubit systems has been studied extensively but mainly in the context of quantum entanglement del Valle (2010); Liao et al. (2011); Fischer and Breuer (2013); Duan et al. (2013); Paneru et al. (2020); Schlosshauer (2019); Aolita et al. (2015); Decordi and Vidiella-Barranco (2017); Bresque et al. (2021); Hewgill et al. (2018); Hu et al. (2018); Karimi et al. (1965); in the present paper, however, the focus is on quantum thermodynamics and the second law of thermodynamics.
@@ II.3 Redfield - Lindblad - Coherences
Then issue of coherences in the steady state and currents is subtle Li et al. (2014); Zhang and Wang (2014); Wichterich et al. (2007). In recent work Fogedby (2022) mentioned in the introduction we presented a detailed discussion of the application of field theoretical methods to open quantum systems. The field theoretical approach is based on a systematic expansion in terms of diagrams and applies a condensed matter quasi-particle approximation in order to implement the Markov limit. The field theoretical method is the basis for the present analysis.
@@ III.1.1 Model
$H=\frac{\omega_{0}}{2}\sigma^{z}$ (22)
@@ III.2.1 Model
$H=\frac{\omega_{1}}{2}\sigma^{z}_{1}+\frac{\omega_{2}}{2}\sigma^{z}_{2}+\lambda(\sigma^{+}_{1}\sigma^{-}_{2}+\sigma^{-}_{1}\sigma^{+}_{2}).$ (56)
@@ III.2.3 Density matrix
Introducing the total kernel $K=K^{A}+K^{B}$, where $K^{A}$ and $K^{B}$ are given by (109) and (114), respectively, the populations $\rho_{nn}$ are determined by the linear equations $\sum_{q}K_{pp,qq}\rho_{qq}=0$. By inspection we note incidentally that $\det K=0$ allowing for a solution. In terms of the parameters $s_{n}=a_{n}+b_{n}$, $n=1\cdots 4$, where $a_{n},b_{n}$ are given (100), we find the normalised populations, i.e., $\sum_{n}\rho_{nn}=1$,
@@ III.2.4 Currents
$q^{A}_{-}=\frac{1}{2}\frac{(\alpha\beta)^{2}g_{A}(\omega_{-})g_{B}(\omega_{-})\omega_{-}}{\beta^{2}g_{A}(\omega_{-})T_{A}+\alpha^{2}g_{B}(\omega_{-})T_{B}}[T_{A}-T_{B}].$ (148)
@@
$q^{A}_{-}=\frac{(\alpha\beta)^{2}g_{A}(\omega_{-})g_{B}(\omega_{-})\omega_{-}}{\beta^{2}g_{A}(\omega_{-})+\alpha^{2}g_{B}(\omega_{-})}[e^{-\omega_{-}/T_{A}}-e^{-\omega_{-}/T_{B}}].$ (152)
@@ III.3 Redfield approach
We have in Fig. 5 depicted the populations $\rho_{nn}$ as a function of the mean temperature $T=(T_{A}+T_{B})/2$ in the range $\Delta T<T<8$ with non equilibrium bias $\Delta T=(T_{A}-T_{B})/2=5.0$ with parameter choice $\omega_{1}=1.0$, $\omega_{2}=2.0$, $g_{A}(\omega_{\pm})=g_{b}(\omega_{\pm})=1.0$, and $\lambda=0.5$. The plot shows that $\rho_{44}$ and $\rho_{44}$ assumes negative value in the range $0<T\lesssim 6.0$, indicating that the Redfield approach is inadequate in analysing the coupled qubit case.
@@ IV Discussion and summary
In this paper we have discussed heat currents in two qubit systems driven by two heat reservoirs maintained at different temperatures in order to test the validity of the second law of thermodynamics in the microscopic domain, an important issue in quantum thermodynamics. The results for a single qubit system driven by two reservoirs are given in Sec. III.1 and are in accordance with the second law. The case of coupled qubits is more involved, the results are presented in Sec. III.2 and are likewise in agreement with the second law. We note that the current expressions in the single qubit and coupled qubit cases have a similar structure. In deriving these results it is essential to apply a global approach and operate in the basis of the full system Hamiltonian. Moreover, in order to avoid an erroneous result it is also essential to use the Lindblad master equation approach and in this way ensure a proper quantum channel with positive populations.